The Surf club at a college in California wants to charter a bus for a surf trip in Mexico. It costs $2600 to rent the bus and each student will pay $370 for lodging, meals, and surfboard rentals.

A) Let C(n) be the total cost for n students going on the trip. Find an equation for C: C(n)=
B) Let M(n) be the mean (or average) cost per student (in dollars per student) if n students go on the trip. Find an equation for M:
M(n)=
C) What is the mean cost if 6 students go? Mean cost =
D) The surf trip will be cancelled unless the mean cost per student is less than $453. What is the minimum number of students needed to go on the trip so it won't be cancelled? Your answer should be an integer.

Answer by Boreal(15235)

C(n)=2600+370(n)
M(n)=C(n)/n=(2600+370(n))/n
Mean for 6 is cost/6 or 2600+2220 divided by 6, 4820/6=$813.33.
453=(2600+370n)/n
453n=2600+370n
83n=2600
n=31.32 and round up to 32 students, where the cost will be 2600+32(370)=14440
divide that by 32 to get $451.25